From GlueXWiki
Jump to: navigation, search

FDC Weekly Meeting

Date: July 30, 2009

Participants: Daniel, Brian, Beni, Mark I., Bill, Simon, Roger, Fernando, Eugene, Rich

Next Meeting: August 6, 2009 @ 1:30 p.m.

General Construction Updates

 - Mark S. has been working exclusively on the solenoid for the last 5 days.
   He should be back on the FDC project starting tomorrow.
 - The parts for the assembly jig on which the detector stack will be
   constructed will be put together as there is time.
 - Bill will take on the preparation for a stack assembly document. A
   draft will be prepared before assembly of the full-scale prototype and
   then fleshed out during the process with relevant photographs taken.
 - The composite frames for the two gas windows have been completed. The
   tensioning of the gas windows and the lamination to the composite rings
   will be done as we have time in the next couple of weeks. 
 - The order for the cathode pigtail cables has been placed as well as the
   order for the connectors to the outside ground lines. There was confusion
   over the status of the conducting epoxy order. It seems that only samples
   have been ordered and there are tests to do to make a final decision.
   We need Mark S. to clear this up.
 - Rich will only be working with us for one more week before going back to 
 - DSC reviewed the status of the Hall B chamber folks wire vendor 
   qualification checks. Apparently the Little Falls Alloys CuBe wire is
   of poor quality compared to the samples from Luma Wire. However, the
   status of our joint wire order with Hall B is unclear. Beni will look
   into this by talking with Steve Christo.
 - One of the dark boxes in EEL 126 will be removed shortly to give us
   some more space. We also talked a bit about longer term space needs.

Wire Frame Update

 - Rich has been working on removing the solder flux from frame P4 and
   is still working on this tedious process. Brian repaired two wires that
   were broken during the cleaning/repair. A number of components have to
   be replaced that (for unknown reasons) were broken off the board. This
   soldering work will take place tomorrow. Rich will then do a final
   cleaning and HV check and the board will be put into storage.
 - The cleaning work that Rich as done on frame P4 has really worked in
   "fixing" the current draw problems that were seen on several busses.
   These arose during out initial cleaning attempt on the boards to remove
   the solder flux. We removed it from the solder pad area but spread it
   around on the board using the Flux-Off material. Hand cleaning is the
   best way. Cleaning the frames in a bath might be the solution in the
   future for contaminated boards. However, the cleaning really needs to be
   done properly by the wire winding folks.
 - Brian will begin wire disconnect repairs on the remaining frame that he
   has not yet worked on. Cleaning, HV checks, and component checks will then 
   be performed. The remaining two wire frames will then need final cleaning
   and component checks before they are certified as done. 
 - We will work to have three wire frames certified and the work on the
   repairs of the fourth frame as we have time.
 - We need to have a dedicated meeting to design what the schedule for
   work on the Phase 3 wire winding will be. The highest priority for
   work scheduling will be doing work on upgrades of the wire winding
   facility. This includes a better strongback design, the design and
   construction of precision combs, modification to the wire winding
   table, and a list of other items with the system. Stay tuned, more
   to come.
 - The wire frame construction document is located at:
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. We will include
   the information on the wire electroplating process shortly.
 - The parts of the jig to use for the wire electroplating operation
   are in our hands. They will be assembled as we find time.
 - Brian received the SEM scans for our wire from the W&M folks. He will
   generate a final QA report with photos and test results for all wires
   when he gets a chance.

Cathode Update

 - Progress updates:
     * The 7th and last cathode was tensioned this past week and was
       laminated to its composite frame.
     * We will modify the layout of our cathode panel alignment fiducials
       to allow us to catch panel rotations. Details were passed on to
     * One of our completed cathode boards is really contaminated with
       epoxy near the two seams. The epoxy manufacturer sent us a solvent 
       that they believe should allow the epoxy to be removed. Testing
       with sample pieces will begin soon.
     * It was found that the cathodes were not sitting flat on the flatness
       system. Bill has come up with a work around and we will begin
       scanning the cathodes next week.
 - Bill stepped us through a work plan for the construction of the two
   cathode sandwiches for the prototype. This work will begin next week
   with the tensioning of the ground planes.
 - Bill is devising a work plan for the prototype construction that he
   will give to Chuck to put into a schedule for tracking.
 - The Fast Electronics group is handling the soldering of the rigid-flex
   assemblies. They should be done within a week. Essentially all of the
   soldering is done, but Fernando was some rigorous QA done on the completed
 - Bill has been working on the cathode construction document building in
   the pictures that are being taken. The current draft of the document 
   is located at /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. 
 - We are standing by waiting for information on the PR for the 2-um board
   material. We have learned that the order has gone out, but the Japanese
   company has still not indicated whether they will fill the order. A
   second company that can produce the board material has been identified.
   Roger was in contact with the U.S. sales representative for the Japanese
   company today.
 - Roger needs to prepare a document for QA/stuffing/cleaning for the
   cathode boards and a similar document for the cathode daughter
   boards and ground boards.

Cooling System Tests

 - Fernando is in the process of writing up the results from his cooling 
   system studies.  Stay tuned for the GlueX note. He indicated that he
   is back to work on this document and it should be done shortly.

Cosmic Ray Test Stand

 - Beni is continuing work to setup the cosmic ray telescope and test
   system in EEL 126. Progress is slow and steady. The gas tubing has been
   laid out and gas is flowing to all chambers.
 - We still have some mysterious behavior with the small-scale prototype
   response with wiggles in the X vs. Y distributions that need to be
   understood. We decided to have a separate meeting to discuss what work
   we would like to do on the small-scale prototype.
 - Beni has started a bench test plan for the full-scale prototype and
   is presently working on it after incorporating feedback. It was also
   suggested to send it to Curtis for feedback. He will post it as a 
   GlueX-note shortly.
 - The parts for the support/rotation of the full-scale prototype in the 
   cosmic ray stand will be assembled as we have time.

Work List

 - The FDC short-term work list has been posted on the FDC web site.  
   This is continually being updated and DSC welcomes any feedback or 
   comments from the group.

Minutes prepared by Daniel. Send any comments or corrections along.